OpenStudy (anonymous):

use the definition of derivatives to compute f(x)= 3x-x^2

OpenStudy (kainui):

Show me your best guess and I'll help.

OpenStudy (anonymous):

Okay thanks.

OpenStudy (anonymous):

\[\frac{ 3(x+h)-(x+h)^{2} -3x-x ^{2}}{ h} \] Is that right so far?

hartnn (hartnn):

just one error -(3x-x^2) will be -3x + x^2

hartnn (hartnn):

after correcting that, you can expand 3(x+h) and (x+h)^2 and notice which terms get cancelled from the numerator.

OpenStudy (anonymous):

yes! I just did. That's why I was confused. So far I have \[\frac{ 3h+2xh+h ^{2} }{ h }\]

hartnn (hartnn):

now what can you factor out from the numerator ?

hartnn (hartnn):

AND, shouldn't that be -2xh ??

hartnn (hartnn):

-(x+h)^2 = -x^2-h^2\(\large -2xh\)

hartnn (hartnn):

and -h^2 too

OpenStudy (anonymous):

ohh I see what I did. so i should always distribute my negatives first?

hartnn (hartnn):

yes, thats preferable

hartnn (hartnn):

now tell me what u get as numerator ?

OpenStudy (anonymous):

so then i get \[3h-2xh-h ^{2}\]and i factor out the h and get 3-2x-h

hartnn (hartnn):

absolutely correct! and since lim h->0 you can now directly plug in h=0 :)

OpenStudy (anonymous):

Yay! Thanks so much for your help. You cleared everything up for me !

OpenStudy (kainui):

Don't forget you can check yourself really easily, and you always should, by taking the derivative the easy way!

hartnn (hartnn):

and whats your final answer ? and yes, cross-checking is always good :)

OpenStudy (anonymous):

3-2x

hartnn (hartnn):

thats correct :) welcome ^_^

OpenStudy (anonymous):

Thanks! :)